can you explain why is 10%

3) Suppose the yen is expected to appreciate by 4% against the pound in one year. If the nominal interest in Japan is 6%, and uncovered interest parity holds, the nominal interest rate in the UK. must be (hint: use the simpler/linear formula) (a) 2%. (b) 5%. (c) 10%.
Answer #1

Forward exchange rate = spot rate x ( 1+ rf) / (1+rd)

rf = foriegn currency investment rate = 6% = 0.06

rd = domestic currency exchange rate = U.K = y

assume

currency value of japan as = X = spot rate

forward exchange rate = X + 4% X = 1.04X

1.04 X = X ( 1+0.06)/(1+y)

1+y = 1.06/1.04

1+y = 1.02

y = 0.02 = 2%

So UK intrest rate = 2%

So UK intrest rate = 2%